OpenStudy (anonymous):

On the graph of the equation 2x − y = 10, what is the value of the y-intercept? -10 -5 5 10

OpenStudy (anonymous):

-10? @satellite73

OpenStudy (anonymous):

i'm just checking my answers

OpenStudy (anonymous):

-10

OpenStudy (anonymous):

2x − y = 10 2x-10= y y= 2x-10 y= mx+c here y intercept = c = -10

OpenStudy (anonymous):

thanks